# Comparative study of the butterfly velocity in holographic QCD models at finite temperature and chemical potential

### v2

@@ 1 Introduction
While chaos has been extensively studied in various classical and quantum systems, exploring its manifestation in Quantum Chromodynamics (QCD) poses unique challenges. Unlike simpler systems, QCD generally operates in a regime where the coupling becomes strong, rendering traditional perturbative methods ineffective. This non-perturbative nature complicates the direct calculation of chaotic indicators such as the Lyapunov exponent and butterfly velocity. Moreover, the confinement of quarks and gluons prevents straightforward observation of chaos signatures in experiments. Recent developments, however, have opened new pathways. The gauge/gravity duality, specifically in the context of holographic QCD, has provided an alternative non-perturbative framework to probe chaotic dynamics in the strongly coupled domain of QCD. In this framework, features of quantum chaos, including out-of-time-order correlators (OTOCs) and pole-skipping phenomena, have been successfully mapped to gravitational dynamics in higher dimensions, offering deep insights into the chaotic properties of QCD-like theories [5, 6]. These developments not only help bridge theoretical gaps but also bring us closer to identifying universal chaotic properties in QCD that could be probed experimentally.
@@
The AdS/CFT correspondence or the gauge-gravity duality [7, 8, 9] is an equivalence between a gravitational theory on a higher-dimensional Anti-de Sitter (AdS) spacetime and a quantum field theory (QFT) living at its boundary in one lower dimension. This duality, originally envisioned in the context of string theory, implies that a strongly coupled quantum field theory, for example, QCD, can be mapped to a classical gravitational theory. This correspondence has been remarkably successful in addressing problems in diverse areas of physics, such as quantum gravity, strongly coupled systems, quark-gluon plasma (QGP), etc, to name just a few. It has provided a novel framework for testing phenomena such as confinement, phase transitions, and quantum chaos in gauge theories, thereby providing an alternate technique to probe QCD behavior in the strong coupling domain and thus non-perturbative QCD. One may refer to [10, 11] for detailed reviews on this topic.
@@
Courtesy of the works [12, 13, 14, 15], the study of chaos in strongly-coupled many-body quantum systems through the lens of gauge/gravity duality has witnessed significant interest. For a comprehensive review of holographic chaos, one may refer to [16]. It was suggested that black holes are the fastest scramblers in nature [17]. In the context of gauge/gravity duality, a thermal state of the boundary CFT corresponds to a black hole in the bulk AdS space. Thus, a rapid thermalization of a local perturbation in the boundary CFT can be understood through the fast scrambling dynamics of the black hole in the dual bulk space. It is also known that black holes in holography can be characterized via quantum chaos [18].
@@
The gauge/gravity duality has also been used extensively to probe chaotic features of QCD and explored in multiple works. In [19, 20], chaos in the hadronic phase of QCD was studied, with a focus laid on homogeneous meson condensates and time-dependent Wilson loops. The investigation of chaos in QCD chiral condensates in $\mathcal{N}=2$ supersymmetric QCD was done in  [21], and the $\mathcal{N}=2$ theory was found to be chaotic. In [22, 23, 24, 25, 26], anisotropic and frame-dependent chaos of a suspended string, corresponding to a quark-antiquark pair in the dual boundary theory, in the holographic deconfined phase was studied. Recently, [27] showed that out-of-time-ordered correlators (OTOC) could also be used to probe the baryonic phase structure in holographic QCD with instantons. Similarly, tools like Krylov complexity and Pole-skipping have also been suggested as order parameters for the deconfinement transition [28, 29]. By now, a lot of work has been done by analyzing open and closed string dynamics to explore chaos in QCD confining and deconfined phases, for instance, see [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
@@
Among the various hallmarks of chaos, the butterfly velocity ($v_{B}$) has emerged as a critical quantity that characterizes the spatial spread of perturbations in quantum many-body systems. Conceptually, $v_{B}$ determines the “speed limit” at which information and chaotic disturbances propagate through a system, echoing the butterfly effect in a spatially extended context. In holographic theories, $v_{B}$ is closely tied to the near black hole horizon dynamics, where it manifests through shockwave geometries and pole-skipping features in retarded correlation functions [45, 46]. Its universality has made it a valuable diagnostic tool in probing the chaotic structure of strongly coupled systems, including QCD-like theories. Notably, $v_{B}$ has been utilized to explore phase transitions and critical points [47], and the anisotropic nature of quark-gluon plasma via holography [48]. More broadly, $v_{B}$ has been related to experimentally relevant observables of QGP: recent work showed that both the heavy-quark drag force and the jet-quenching parameter in a thermal plasma are directly controlled by the butterfly velocity [49]. Such findings suggest that chaotic dynamics may manifest in transport and energy-loss phenomena in strongly coupled QCD matter. This makes it not only a theoretical construct but also a potential observable, connecting chaotic dynamics with experimentally measurable transport phenomena.
@@
In this work, our main aim is to investigate the butterfly velocity in the deconfined phase of QCD using holography and study the influence of finite chemical potential $\mu$ and temperature $T$ on QCD chaotic dynamics. These parameters are essential variables in the QCD phase diagram and qualitatively alter its behavior, and we particularly aim to study their effects on the butterfly velocity in various top-down and bottom-up holographic QCD models. The top-down models, while firmly grounded and can be constructed from higher-dimensional string theory and thus are more robust regarding the validity of the duality, often fail to capture key features of QCD. In contrast, bottom-up phenomenological models, typically constructed in an ad hoc manner with limited string-theoretic justification, are more successful in reproducing many desirable aspects of QCD. To make our analysis more concrete and thorough, we consider both top-down and bottom-up holographic QCD models. This includes top-down 1RCBH and 2RCBH holographic models [50, 51], potential reconstruction based analytic bottom-up model [52, 53], and a numerical bottom-up model [54, 55]. All these models are based on the Einstein-Maxwell-dilaton gravity action, distinguished by distinct forms of the dilaton potential and the gauge-kinetic coupling function, unique in their own rights, and have been thoroughly used in the AdS/QCD literature to probe and explore various QCD-related features from holography.
@@
For each of these holographic models, we compute the butterfly velocity using three distinct methods—entanglement wedge reconstruction, out-of-time-ordered correlators (OTOCs), and pole-skipping—and find that all three yield the same result. Note that such an equivalence between these methods has been previously shown in higher order curvature gravity theories [56], Lifshitz theories [57], hyperscaling-violating Lifshitz theories [58], $T\bar{T}$ deformed setup [59], etc., among various other holographic theories. For other related works, see [60, 61, 62, 63, 64, 65]. Our work, therefore, establishes the equivalence between these three methods in AdS/QCD models. Note that, the conceptual explanation for the equivalence between the three methods was presented recently in [66]. Importantly, for the first three models, we obtain analytic results for the butterfly velocity, whereas for the fourth model, we obtain results numerically. We further analyze the thermal and chemical dependent profile of the butterfly velocity and find that it exhibits similar features across models. In particular, it increases and then saturates to conformal values as the temperature increases, while it decreases with the chemical potential, suggesting an intricate interplay between temperature, chemical potential, and the butterfly velocity.
@@
The structure of the paper is as follows. In section 2, we discuss in detail the three methods, namely entanglement wedge, OTOC, and pole-skipping, to compute butterfly velocity $v_{B}$. We consider a very general form of planar and isotropic metric and arrive at a general expression for $v_{B}$ using each of the three methods. In section 3, we compute the butterfly velocity in 1RCBH and 2RCBH holographic models. In sections 4 and 5, we repeat the butterfly computation in bottom-up holographic QCD models of [52] and [55], respectively. Finally, in section 6, we conclude and discuss future directions to extend your work.
@@ 2.1 Entanglement wedge
In [67], the delocalization dynamics of a local operator in a thermal equilibrium state were analyzed through the lens of holographic duality. When a local operator acts on the boundary thermal state, it introduces a perturbation represented in the bulk as a particle propagating toward the horizon of a static black hole. Over time, the spatial support of the operator in the boundary theory expands, requiring increasingly larger boundary subregions to reconstruct information about the perturbation. This growth is governed by the butterfly velocity $v_{B}$, which characterizes the effective light-cone speed of operator spreading in chaotic systems. In the bulk description, the smallest boundary subregion capable of reconstructing the perturbation corresponds to an entanglement wedge—the bulk region bounded by the Hubeny-Rangamani-Takayanagi (HRT) surface anchored to the subregion. As the particle falls inward, the radial depth of its trajectory determines the minimal boundary subregion whose entanglement wedge contains the particle. This establishes a dynamical relationship: the time-dependent spatial extent of the operator on the boundary is dual to the radial infall of the particle in bulk, with $v_{B}$ dictating the rate at which the required boundary subregion expands. Figure 1 illustrates this process, where the entanglement wedge (yellow region) associated with a boundary subregion grows to encompass the infalling perturbation (red dot) at successive times.
@@
This mechanism provides a gravitational explanation for entanglement wedge subregion duality in AdS/CFT, demonstrating how the emergent light-cone structure of chaotic systems, encoded in $v_{B}$, directly governs the holographic reconstruction of the bulk locality from boundary data. The saturation of $v_{B}$-dependent bounds on operator growth further underscores the connection between entanglement dynamics, scrambling, and the geometric emergence of causal structure in holographic theories [14, 72]. To compute the entanglement wedge, we start with the planar black hole with the following general form of the metric
@@
here $g(z)$ is the blackening function and $z$ is the usual radial coordinate. [^1] Throughout this paper, we will denote the holographic radial coordinate by $z$ instead of $r$. Accordingly, the $z$ coordinate runs from $z=z_{h}$ (horizon) to $z=\infty$ (asymptotic boundary). We reserve the coordinate $r$ to denote the radial coordinate in the $x^{i}$ plane. The horizon is defined by the condition $g(z_{h})=0$, with $z_{h}$ being the horizon radius. Note that the spatial part of the metric is isotropic and planar. Since the position of the particle is time-dependent and not the background geometry itself, the Ryu-Takayanagi (RT) surface [73] is taken as the extremal surface in this case and not the Hubeny-Rangamani-Takayanagi (HRT) surface [74]. Thus, for our purposes, the entanglement wedge is a $t=\text{constant}$ hypersurface bounded by the RT surface and the boundary subregion.
The entanglement entropy or the area functional is given as follows
@@
where $\xi$ are the appropriate coordinates on the codimension-two surface in the bulk, and $\gamma$ is the determinant of the induced metric on the surface. Extremizing $S_{EE}$ (or equivalently, the induced metric $\gamma$) gives rise to the equation of the RT surface. Since the surface is embedded in a $t=\text{constant}$ hypersurface, the $tt$ component of the induced metric vanishes. Further, we take $r$ as the radial coordinate in the $x^{i}$ direction ($r=|x|$) and parametrize $z$ with $r$, i.e., $z=z(r)$. Hence, from (2.1) we get the following induced metric
@@
Next, analogous to the computation of [67], we do a near-horizon analysis of the RT surface to compute the butterfly velocity. Near the horizon, we can take the following form of $z(r)$
@@
here the subscript $o$ indicates that the functions are evaluated at $z=z_{h}$ and the prime indicates the derivative with respect to $z$. Extremizing (2.5), we arrive at the following differential equation for $u(r)$
@@
The RT surface starts to depart the near-horizon region for large $r$. To put it precisely, the RT surface stays close to the horizon up to the point where $\epsilon u(r)^{2}\sim$ $\mathcal{O}(\frac{1}{z_{h}})$, and after that, the surface departs the near-horizon region and reaches the boundary to order one distance. As suggested in [56], we can encapsulate this behavior via the following Ansätze
@@
where $n$ is some positive integer. Also, the particle touches the tip of the RT surface at all times. Thus, taking the tip of the RT surface to be the origin, we can set $u(o,t)\sim e^{-2\pi t/\beta}$. Thus $u(r,t)$ is given as follows
@@
The rate at which the particle propagates towards the horizon is the rate at which the size of the applied operator grows (we make the notion of “size” more precise in section 2.2). Thus, from (2.8), the butterfly velocity is given as
@@
To determine $v_{B}$, we need to determine $\sigma$. To determine $\sigma$, we substitute the Ansätze (2.7) into (2.6). Dropping higher-order terms in $1/r$, we get the following expression for $\sigma$
@@
It is worth noting that if the temperature of the black hole is not a function of the chemical potential (like in certain holographic models), the temperature does not become zero for any value of the chemical potential. Also, the function $f(z)$ cannot vanish at the horizon because $g_{o}=0$ (from (2.1)). This implies that $v_{B}$ in such models is never exactly zero for any finite value of the chemical potential. We will consider such an example in section 3.
@@
2.2 Shock Wave and Out-of-time-ordered-correlators (OTOCs)
We are interested in the spreading of the quantum butterfly effect in spatial directions of the boundary QCD systems at strong couplings. The same is intrinsically related to the size of perturbative local operators and can be obtained from the OTOCs of local operators at the boundary. In the dual gravity picture, one can think of this as injecting a small number of quanta, which propagate toward the black hole in the bulk and create a localized shockwave. The spatial region where the shockwave has non-trivial support determines the size of the corresponding boundary operator. Accordingly, one can evaluate the butterfly velocity by computing how this special region grows by sending perturbations at progressively earlier times.
The operator size or the spreading of the quantum butterfly effect in a system can be quantified via the following commutator

### v3

@@ 1 Introduction
While chaos has been extensively studied in various classical and quantum systems, exploring its manifestation in quantum chromodynamics (QCD) poses unique challenges. Unlike simpler systems, QCD generally operates in a regime where the coupling becomes strong, rendering traditional perturbative methods ineffective. This nonperturbative nature complicates the direct calculation of chaotic indicators such as the Lyapunov exponent and butterfly velocity. Moreover, the confinement of quarks and gluons prevents straightforward observation of chaos signatures in experiments. Recent developments, however, have opened new pathways. The gauge/gravity duality, specifically in the context of holographic QCD, has provided an alternative nonperturbative framework to probe chaotic dynamics in the strongly coupled domain of QCD. In this framework, features of quantum chaos, including out-of-time-order correlators (OTOCs) and pole-skipping phenomena, have been successfully mapped to gravitational dynamics in higher dimensions, offering deep insights into the chaotic properties of QCD-like theories [5, 6]. These developments not only help bridge theoretical gaps but also bring us closer to identifying universal chaotic properties in QCD that could be probed experimentally.
@@
The AdS/CFT correspondence or the gauge-gravity duality [7, 8, 9] is an equivalence between a gravitational theory on a higher-dimensional Anti-de Sitter (AdS) spacetime and a quantum field theory (QFT) living at its boundary in one lower dimension. This duality, originally envisioned in the context of string theory, implies that a strongly coupled quantum field theory, for example, QCD, can be mapped to a classical gravitational theory. This correspondence has been remarkably successful in addressing problems in diverse areas of physics, such as quantum gravity, strongly coupled systems, quark-gluon plasma (QGP), etc, to name just a few. It has provided a novel framework for testing phenomena such as confinement, phase transitions, and quantum chaos in gauge theories, thereby providing an alternate technique to probe QCD behavior in the strong coupling domain and thus nonperturbative QCD. One may refer to [10, 11] for detailed reviews on this topic.
@@
Courtesy of the works [12, 13, 14, 15], the study of chaos in strongly coupled many-body quantum systems through the lens of gauge/gravity duality has witnessed significant interest. For a comprehensive review of holographic chaos, one may refer to [16]. It was suggested that black holes are the fastest scramblers in nature [17]. In the context of gauge/gravity duality, a thermal state of the boundary conformal field theory (CFT) corresponds to a black hole in the bulk AdS space. Thus, a rapid thermalization of a local perturbation in the boundary CFT can be understood through the fast scrambling dynamics of the black hole in the dual bulk space. It is also known that black holes in holography can be characterized via quantum chaos [18].
@@
The gauge/gravity duality has also been used extensively to probe chaotic features of QCD and explored in multiple works. In [19, 20], chaos in the hadronic phase of QCD was studied, with a focus laid on homogeneous meson condensates and time-dependent Wilson loops. The investigation of chaos in QCD chiral condensates in $\mathcal{N}=2$ supersymmetric QCD was done in  [21], and the $\mathcal{N}=2$ theory was found to be chaotic. In [22, 23, 24, 25, 26], anisotropic and frame-dependent chaos of a suspended string, corresponding to a quark-antiquark pair in the dual boundary theory, in the holographic deconfined phase was studied. Recently, [27] showed that out-of-time-ordered correlators (OTOC) could also be used to probe the baryonic phase structure in holographic QCD with instantons. Similarly, tools like Krylov complexity and pole-skipping have also been suggested as order parameters for the deconfinement transition [28, 29]. By now, a lot of work has been done by analyzing open and closed string dynamics to explore chaos in QCD confining and deconfined phases, for instance, see [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41].
@@
Among the various hallmarks of chaos, the butterfly velocity ($v_{B}$) has emerged as a critical quantity that characterizes the spatial spread of perturbations in quantum many-body systems. Conceptually, $v_{B}$ determines the “speed limit” at which information and chaotic disturbances propagate through a system, echoing the butterfly effect in a spatially extended context. In holographic theories, $v_{B}$ is closely tied to the near black hole horizon dynamics, where it manifests through shockwave geometries and pole-skipping features in retarded correlation functions [46, 47]. Its universality has made it a valuable diagnostic tool in probing the chaotic structure of strongly coupled systems, including QCD-like theories. Notably, $v_{B}$ has been utilized to explore phase transitions and critical points [48] and the anisotropic nature of quark-gluon plasma via holography [49]. More broadly, $v_{B}$ has been related to experimentally relevant observables of QGP: recent work showed that both the heavy-quark drag force and the jet-quenching parameter in a thermal plasma are directly controlled by the butterfly velocity [50]. Such findings suggest that chaotic dynamics may manifest in transport and energy-loss phenomena in strongly coupled QCD matter. This makes it not only a theoretical construct but also a potential observable, connecting chaotic dynamics with experimentally measurable transport phenomena.
@@
In this work, our main aim is to investigate the butterfly velocity in the deconfined phase of QCD using holography and study the influence of finite chemical potential $\mu$ and temperature $T$ on QCD chaotic dynamics. These parameters are essential variables in the QCD phase diagram and qualitatively alter its behavior, and we particularly aim to study their effects on the butterfly velocity in various top-down and bottom-up holographic QCD models. The top-down models, while firmly grounded and can be constructed from higher-dimensional string theory and thus are more robust regarding the validity of the duality, often fail to capture key features of QCD. In contrast, bottom-up phenomenological models, typically constructed in an ad hoc manner with limited string-theoretic justification, are more successful in reproducing many desirable aspects of QCD. To make our analysis more concrete and thorough, we consider both top-down and bottom-up holographic QCD models. This includes top-down 1 R-charge black hole (1RCBH) and 2RCBH holographic models [51, 52], potential reconstruction-based analytic bottom-up model [53, 54], and a numerical bottom-up model [55, 56]. All these models are based on the Einstein-Maxwell-dilaton gravity action, distinguished by distinct forms of the dilaton potential and the gauge-kinetic coupling function, unique in their own rights, and have been thoroughly used in the AdS/QCD literature to probe and explore various QCD-related features from holography.
@@
For each of these holographic models, we compute the butterfly velocity using three distinct methods — entanglement wedge reconstruction, OTOCs, and pole skipping — and find that all three yield the same result. Note that such an equivalence between these methods has been previously shown in higher-order curvature gravity theories [57], Lifshitz theories [58], hyperscaling-violating Lifshitz theories [59], $T\bar{T}$ deformed setup [60], etc., among various other holographic theories. For other related works, see [61, 62, 63, 64, 65, 66]. Our work, therefore, establishes the equivalence between these three methods in AdS/QCD models. Note that the conceptual explanation for the equivalence between the three methods was presented recently in [67]. Importantly, for the first three models, we obtain analytic results for the butterfly velocity, whereas for the fourth model, we obtain results numerically. We further analyze the thermal- and chemical-dependent profile of the butterfly velocity and find that it exhibits similar features across models. In particular, it increases and then saturates to conformal values as the temperature increases, while it decreases with the chemical potential, suggesting an intricate interplay between temperature, chemical potential, and the butterfly velocity.
@@
The structure of the paper is as follows. In Section 2, we discuss in detail the three methods, namely entanglement wedge, OTOC, and pole skipping, to compute butterfly velocity $v_{B}$. We consider a very general form of planar and isotropic metric and arrive at a general expression for $v_{B}$ using each of the three methods. In Section 3, we compute the butterfly velocity in 1RCBH and 2RCBH holographic models. In Sections 4 and 5, we repeat the butterfly computation in bottom-up holographic QCD models of [53] and [56], respectively. Finally, in Section 6, we conclude and discuss future directions to extend our work.
@@ 2.1 Entanglement wedge
In [68], the delocalization dynamics of a local operator in a thermal equilibrium state were analyzed through the lens of holographic duality. When a local operator acts on the boundary thermal state, it introduces a perturbation represented in the bulk as a particle propagating toward the horizon of a static black hole. Over time, the spatial support of the operator in the boundary theory expands, requiring increasingly larger boundary subregions to reconstruct information about the perturbation. This growth is governed by the butterfly velocity $v_{B}$, which characterizes the effective light-cone speed of operator spreading in chaotic systems. In the bulk description, the smallest boundary subregion capable of reconstructing the perturbation corresponds to an entanglement wedge – the bulk region bounded by the Hubeny-Rangamani-Takayanagi (HRT) surface anchored to the subregion. As the particle falls inward, the radial depth of its trajectory determines the minimal boundary subregion whose entanglement wedge contains the particle. This establishes a dynamical relationship: the time-dependent spatial extent of the operator on the boundary is dual to the radial infall of the particle in the bulk, with $v_{B}$ dictating the rate at which the required boundary subregion expands. Figure 1 illustrates this process, where the entanglement wedge (yellow region) associated with a boundary subregion grows to encompass the infalling perturbation (red dot) at successive times.
@@
This mechanism provides a gravitational explanation for entanglement wedge subregion duality in AdS/CFT, demonstrating how the emergent light-cone structure of chaotic systems, encoded in $v_{B}$, directly governs the holographic reconstruction of the bulk locality from boundary data. The saturation of $v_{B}$-dependent bounds on operator growth further underscores the connection between entanglement dynamics, scrambling, and the geometric emergence of causal structure in holographic theories [14, 73]. To compute the entanglement wedge, we start with the planar black hole with the following general form of the metric:
@@
here $g(z)$ is the blackening function and $z$ is the usual radial coordinate. [^1] Throughout this paper, we will denote the holographic radial coordinate by $z$ instead of $r$. Accordingly, the $z$ coordinate runs from $z=z_{h}$ (horizon) to $z=\infty$ (asymptotic boundary). We reserve the coordinate $r$ to denote the radial coordinate in the $x^{i}$ plane. The horizon is defined by the condition $g(z_{h})=0$, with $z_{h}$ being the horizon radius. Note that the spatial part of the metric is isotropic and planar. Since the position of the particle is time dependent and not the background geometry itself, the Ryu-Takayanagi (RT) surface [74] is taken as the extremal surface in this case and not the HRT surface [75]. Thus, for our purposes, the entanglement wedge is a $t=\text{constant}$ hypersurface bounded by the RT surface and the boundary subregion.
The entanglement entropy or the area functional is given as follows:
@@
where $\xi$ are the appropriate coordinates on the codimension-two surface in the bulk, and $\gamma$ is the determinant of the induced metric on the surface. Extremizing $S_{EE}$ (or, equivalently, the induced metric $\gamma$) gives rise to the equation of the RT surface. Since the surface is embedded in a $t=\text{constant}$ hypersurface, the $tt$ component of the induced metric vanishes. Further, we take $r$ as the radial coordinate in the $x^{i}$ direction ($r=|x|$) and parametrize $z$ with $r$, i.e., $z=z(r)$. Hence, from (2.1) we get the following induced metric:
@@
Next, analogous to the computation of [68], we do a near-horizon analysis of the RT surface to compute the butterfly velocity. Near the horizon, we can take the following form of $z(r)$:
@@
where the subscript $o$ indicates that the functions are evaluated at $z=z_{h}$ and the prime indicates the derivative with respect to $z$. Extremizing (2.5), we arrive at the following differential equation for $u(r)$:
@@
The RT surface starts to depart the near-horizon region for large $r$. To put it precisely, the RT surface stays close to the horizon up to the point where $\epsilon u(r)^{2}\sim$ $\mathcal{O}(\frac{1}{z_{h}})$, and after that, the surface departs the near-horizon region and reaches the boundary to order one distance. As suggested in [57], we can encapsulate this behavior via the following ansatz:
@@
where $n$ is some positive integer. Also, the particle touches the tip of the RT surface at all times. Thus, taking the tip of the RT surface to be the origin, we can set $u(o,t)\sim e^{-2\pi t/\beta}$. Thus $u(r,t)$ is given as follows:
@@
The rate at which the particle propagates toward the horizon is the rate at which the size of the applied operator grows (we make the notion of “size” more precisely in Section 2.2). Thus, from (2.8), the butterfly velocity is given as
@@
To determine $v_{B}$, we need to determine $\sigma$. To determine $\sigma$, we substitute the ansatz (2.7) into (2.6). Dropping higher-order terms in $1/r$, we get the following expression for $\sigma$:
@@
It is worth noting that, if the temperature of the black hole is not a function of the chemical potential (like in certain holographic models), the temperature does not become zero for any value of the chemical potential. Also, the function $f(z)$ cannot vanish at the horizon because $g_{o}=0$ [from (2.1)]. This implies that $v_{B}$ in such models is never exactly zero for any finite value of the chemical potential. We will consider such an example in Section 3.
@@
2.2 Shock wave and OTOCs
We are interested in the spreading of the quantum butterfly effect in spatial directions of the boundary QCD systems at strong couplings. The same is intrinsically related to the size of perturbative local operators and can be obtained from the OTOCs of local operators at the boundary. In the dual gravity picture, one can think of this as injecting a small number of quanta, which propagate toward the black hole in the bulk and create a localized shock wave. The spatial region where the shock wave has nontrivial support determines the size of the corresponding boundary operator. Accordingly, one can evaluate the butterfly velocity by computing how this special region grows by sending perturbations at progressively earlier times.
The operator size or the spreading of the quantum butterfly effect in a system can be quantified via the following commutator: